OpenStudy (anonymous):

Find dy/dx : y=sin^-1[ x{ 1-x }^1/2 - { x }^1/2 {1-( x )^2}^1/2]

OpenStudy (john_es):

I'll rewrite it, \[y=\arcsin (x\sqrt{1-x}-\sqrt{x}\sqrt{1-x^2})\] is this the y you wrote?

OpenStudy (john_es):

Is this the case, then, the derivative is, \[dy/dx=\frac{-\frac{\sqrt{1-x^2}}{2 \sqrt{x}}+\frac{x^{3/2}}{\sqrt{1-x^2}}-\frac{x}{2 \sqrt{1-x}}+\sqrt{1-x}}{\sqrt{1-\left(\sqrt{1-x} x-\sqrt{x} \sqrt{1-x^2}\right)^2}}\]If you need step by step, tell me.

OpenStudy (anonymous):

Yes,please!!

OpenStudy (john_es):

You need to apply the derivative of the arcsin, and chain rule \[(\arcsin (f(x)))'=\frac{f'(x)}{\sqrt{1-(f(x))^2}}\] And of course, the derivative of the square root, \[(\sqrt{f(x)})'=\frac{f'(x)}{2\sqrt{f(x)}}\] With these tools, you can follow the process.

OpenStudy (anonymous):

Got it!!

OpenStudy (john_es):

Ok. ;)

OpenStudy (anonymous):

And thanks.. ;)

OpenStudy (john_es):

You're welcome.

OpenStudy (anonymous):

One more question!!! Is there a way of substituting some value in place of x and solving the problem ?

OpenStudy (john_es):

Yes, of course, with the final formula of the derivative, plug the value you want there and you will have the solution. Although, as you can see, it will be a little clumsy to plug in the value in this big formula.

OpenStudy (anonymous):

Hmmm..I see...